OpenStudy (anonymous):

find the sum of the following infinite geometric series, if it exists. 1/2 + (-1/4)+(1/8)+(-1/16)+...

OpenStudy (anonymous):

\[\frac{\frac 12 }{1+\frac 12}= \frac 13 \]

OpenStudy (anonymous):

Try first by grouping terms as follows. \[\left( \frac{1}{2}-\frac{1}{4}\right) + \left(\frac{1}{8}-\frac{1}{16}\right)+\cdots=\frac{1}{4}+\frac{1}{16}+\cdots=\sum_{n=1}^\infty \left(\frac{1}{4} \right)^n\]This is now just a simple geometric series.

OpenStudy (campbell_st):

find the common ratio by comparing the terms \[\frac{T _{2}}{T _{1}} = \frac{T _{3}}{T2} = \frac{T _{4}}{T _{3}}...\] if |r|<1 use \[s=\frac{a}{1-r}\] a = 1st term r = common ratio

OpenStudy (anonymous):

First Number a: 1/2 Ratio: -1/2 Formula of Finding Sum of Infinite Geometric Series when |r|<1 is \[{a}/{1-r}\] (1/2)/(3/2)=1/3

OpenStudy (anonymous):

okay thanks I think i get it :)